CH^{+} AND CH EMISSION FROM THE RED RECTANGLE

LEWIS M. HOBBS, JULIE A. THORBURN, DONALD G. YORK, \underline{TAKESHI OKA}, Department of Astronomy and Astrophysics, The University of Chicago, Chicago IL 60637; THEODORE P. SNOW, Center for Astrophysics and Space Astronomy University of Colorado, Boulder, CO 80309-0389; and JOHN BARENTINE, New Mexico State University, Apache Point Observatory, Sunspot, New Mexico, 88349-0059.

Using the 3.5 meter telescope at the Apache Point Observatory, we have observed the 0 - 0, 1 - 0, 0 - 1, 1 - 2, and 2 - 1 bands of the Ã^{1}Π → X̃^{1}Σ^{+} transition (origin at 4238 Å) of CH^{+} and the 0 - 0 band of the Ã^{2}Δ → X̃^{2}Π transition (4307 Å) of CH in emission from HD 44179, the central star of the Red Rectangle. The 0 - 0 band emission of CH^{+} was reported earlier,^{a} but all other emission bands are new. The CH^{+} emission lines are observed up to the Q(7) and the P (6) lines demonstrating population of CH^{+} up to at least J = 6 in the ground state. Since CH^{+} has a large dipole moment (1.656 D,^{b} 1.804 D^{c}) the spontaneous emission of the J = 6 → 5 rotational transition at 165.96 cm^{−1} is fast (the Einstein coefficient A = 1.81 s^{−1}, 2.15 s^{−1}). In order to populate up to the J = 6 level, CH^{+} must be in an environment with a high density on the order of 10^{9} cm^{−3}. It has to be a cloud of H atoms since H_{2} will immediately react with CH^{+} to produce CH_{2}^{+}. Existence of such a high density H I cloud is remarkable. The observed CH emission, on the other hand, is limited to the R_{2}(1/2), R_{1}(3/2) and R_{1}(5/2) lines all of which can be reached from the lowest two levels in the ground state that are separated by only 17.9 cm^{−1} and spontaneous emission between them is slow. This clearly demonstrates that CH^{+} and CH exist in entirely different regions of the cloud surrounding HD 44179. A more detailed analysis will be given.

aS. P. Balm and M. Jura, A&A 261 L25 (1992)

bF. R. Ornellas and F. B. C. Machado, J. Chem. Phys. 84 1296 (1986)

cH. Sun and K. F. Freed, J. Chem. Phys. 88 2659 (1988)